OpenStudy (anonymous):

to turn 11/25 into a decimal, I know the next step is going to be 25 divided by 11.00, which results into 2.272, but according to the book answer it says that the result is 0.44?? Can someone help me? thanks!

OpenStudy (shane_b):

That's because 11/25 is 11 divided by 25....NOT 25 divided by 11 :P

OpenStudy (anonymous):

Ohhhhhhhhhhhhhhhhhh!! ... Brain fart!! I need to take a break ...

OpenStudy (anonymous):

\[\frac{11}{25}=\frac{11}{25}\times \frac{4}{4}=\frac{44}{100}=.44\]

OpenStudy (anonymous):

thanks satellite73

OpenStudy (anonymous):

yw